The following table presents the heights (in inches) of a sample of college basketball players.


Height(in) Frequency

68-71 11

72-75 56

76-79 54

80-83 40

84-87 13

Considering the data to be a population. Please find each of these measures of central tendency and measure of variation:

i. Mean

ii. Median

iii. Mode

iv. Midrange

v. Range

vi. Variance

vii. Standard deviation

viii. Range Rule of Thumb

ix. Chebyshev’s Theorem


1
Expert's answer
2021-12-06T16:53:11-0500

i.

Mean = "\frac{\sum m_if_i}{n}=77.22"

where mi is midpoint of class,

fi is frequency of class


ii)

Median = "L+\frac{n/2-cf}{f}c=75.5+4\frac{87-67}{54}=76.98"


The median class is 75.5-79.5

L=lower boundary point of median class =75.5

n=Total frequency =174

cf=Cumulative frequency of the class preceding the median class =67

f=Frequency of the median class =54

c=class length of median class =4


iii)

Mode = "L+\frac{f_1-f_0}{2f_1-f_0-f_2}c=71.5+4\frac{56-11}{2\cdot 56-11-54}=75.33"


 The mode class is 71.5-75.5.

L=lower boundary point of mode class =71.5

f1= frequency of the mode class =56

f0= frequency of the preceding class =11

f2= frequency of the succedding class =54

c= class length of mode class =4


iv)

Midrange = "\frac{68+87}{2}=77.5"


v)

Range = "87-68=19"


vi)

Variance = "\sigma^2=\frac{\sum m^2_if_i-\frac{(\sum m_if_i)^2}{n}}{n}=17.58"


vii)

Standard deviation = "\sigma=\sqrt{\frac{\sum m^2_if_i-\frac{(\sum m_if_i)^2}{n}}{n}}=4.19"


viii)

standard deviation:

"\sigma=R/4=19/4=4.75"

where R is range


ix)

at least 3/4 of the data lie within two standard deviations of the mean, that is, in the interval:

"\mu\pm 2\sigma=77.22\pm 2\cdot4.19=(68.84,85.6)"

at least 8/9 of the data lie within three standard deviations of the mean, that is, in the interval:

"\mu\pm 3\sigma=77.22\pm 3\cdot4.19=(64.65,89.79)"

at least 1-1/k2 of the data lie within k standard deviations of the mean, that is, in the interval:

"\mu\pm k\sigma"